Find x in the solution of the system 3x + y = 2 and 2x - 3y = 16.
Thanks

Answer by fractalier(6550)

To find x, we will eliminate y by multiplying the first equation by 3 and adding it to the second equation...
3x + y = 2
2x - 3y = 16
we get
2x - 3y = 16
9x + 3y = 6
and
11x = 22
x = 2